Suppose in a particular labor market, the demand for labor is given by the equation LD = 120 – 3W and that the labor supply in this market for native-born citizens is given by LN = 3W, while the supply curve of immigrants in this market is given by LI = 2W, where L represents the number of workers, W is the wage expressed in real terms.
Answer #1

Total labor supply= labor supply + immigrant labor

total labor supply=Ls=5W

and Labor Demand=120-3W

equilibrium is achieved when Ls=Ld

120-3W=5W

wage=120/8=15 and total labor=75

native born labor supply=15*3=45

and immigrant =15*2=30